# ANLS* - A Universal Document Processing Metric for Generative Large Language Models

### v4

@@ Appendix A Results GPT
Dataset Method gpt-3.5-turbo-16k-0613 gpt-3.5-turbo-0125 gpt-4-1106-preview gpt-4-turbo-2024-04-09 gpt-4-vision-preview-1106
DocVQA Simple 0.576 0.550 0.607 0.589 0.759
Latin Prompting 0.620 0.626 0.699 0.631 -
SFT (Ours) 0.768 0.733 0.790 0.770 -
MPDocVQA Simple 0.507 0.461 0.635 0.606 0.708
Latin Prompting 0.489 0.507 0.739 0.656 -
SFT (Ours) 0.724 0.731 0.781 0.760 -
Kleister Charity Simple 0.527 0.564 0.743 0.708 0.751
Latin Prompting 0.477 0.493 0.735 0.728 -
SFT (Ours) 0.534 0.548 0.763 0.733 -
Kleister NDA Simple 0.361 0.575 0.695 0.633 0.664
Latin Prompting 0.412 0.593 0.705 0.643 -
SFT (Ours) 0.315 0.562 0.703 0.659 -
SROIE Simple 0.509 0.706 0.835 0.849 0.834
Latin Prompting 0.723 0.708 0.851 0.847 -
SFT (Ours) 0.792 0.649 0.873 0.890 -
VRDU AD Buy Simple 0.414 0.084 0.553 0.483 0.640
Latin Prompting 0.424 0.211 0.586 0.506 -
SFT (Ours) 0.676 0.452 0.770 0.712 -
VRDU Registration Simple 0.586 0.615 0.676 0.609 0.665
Latin Prompting 0.606 0.630 0.673 0.614 -
SFT (Ours) 0.637 0.654 0.711 0.590 -

### v5

@@ Appendix A Results GPT
Dataset Method gpt-3.5-turbo-16k-0613 gpt-3.5-turbo-0125 gpt-4-1106-preview gpt-4-turbo-2024-04-09 gpt-4-vision-preview-1106 gpt-4o-2024-05-13
DocVQA Simple 0.576 0.550 0.607 0.589 0.759 0.656
Latin Prompting 0.620 0.626 0.699 0.631 - 0.718
SFT (Ours) 0.768 0.733 0.790 0.770 - 0.747
MPDocVQA Simple 0.507 0.461 0.635 0.606 0.708 0.686
Latin Prompting 0.489 0.507 0.739 0.656 - 0.718
SFT (Ours) 0.724 0.731 0.781 0.760 - 0.762
Kleister Charity Simple 0.527 0.564 0.743 0.708 0.751 0.711
Latin Prompting 0.477 0.493 0.735 0.728 - 0.721
SFT (Ours) 0.534 0.548 0.763 0.733 - 0.737
Kleister NDA Simple 0.361 0.575 0.695 0.633 0.664 0.649
Latin Prompting 0.412 0.593 0.705 0.643 - 0.646
SFT (Ours) 0.315 0.562 0.703 0.659 - 0.655
SROIE Simple 0.509 0.706 0.835 0.849 0.834 0.887
Latin Prompting 0.723 0.708 0.851 0.847 - 0.863
SFT (Ours) 0.792 0.649 0.873 0.890 - 0.897
VRDU AD Buy Simple 0.414 0.084 0.553 0.483 0.640 0.596
Latin Prompting 0.424 0.211 0.586 0.506 - 0.655
SFT (Ours) 0.676 0.452 0.770 0.712 - 0.753
VRDU Registration Simple 0.586 0.615 0.676 0.609 0.665 0.605
Latin Prompting 0.606 0.630 0.673 0.614 - 0.611
SFT (Ours) 0.637 0.654 0.711 0.590 - 0.605